| 12/20/2004 | |||
| 0:34 | ESC turned off the flight computer | JH | |
| 0:36 | ESC turned on the flight computer | JH | |
| They said that they are getting something. | |||
| ESC tried to reconnect with cdps2 | |||
| 0:48 | restarted the cdaq-cream program for data flow | JH | |
| 0:51 | The connection was restored | SO | |
| 1:00 | We started the processing that parameters were turned on. | SO | |
| but SCD looks like power off | |||
| 1:03 | sent command : SCI_CAL_HV_LEVEL 0x05,0 ==>CAL MB5 turned off | YS | |
| Af first, I send "hv_on_hds_1220.stol", then bias went up but HV doesn’t goes up. Calorimeter's HV still goes up. | |||
| 1:20 | When we have commanding window again, we sent HV on commands for hds. | YS | |
| Set HV for cal to 6KV (2000) | |||
| 1:28 | power on SCD | JH | |
| ESC said they need to switch CDM flight computer. They said it will take 10 minues | |||
| 1:37 | While ramping up for HDS, HV of HDS for H11 and H12 fall down to zero. It looked like tripped. | ||
| 1:51 | HV on SCD | JH | |
| 1:59 | For tripped H11 AND h12 | YS | |
| Send "SCI_HV_LEVEL 0x17, 0" // set HV 0 for H11 | |||
| Send "SCI_HV_LEVEL 0x18, 0" //set HV 0 for H12 | |||
| send bias off "SCI_HDS_BIAS 0x17, OFF" | |||
| send bias off "SCI_HDS_BIAS 0x18, OFF" | |||
| Send Power off "SCI_HDS_DC_POWER 0x0D, OFF" | |||
| Send Power off "SCI_HDS_DC_POWER 0x0E, OFF" | |||
| Click "SYNC HV LEVEL" | |||
| POWER ON H11 and H12 | |||
| BIAS on H11 and H12 | |||
| Started ramping up HV to 6kV level (2000). | |||
| But, it tripped again around 3kV. | |||
| I'm going to leave them at 0 HV | |||
| Send "SCI_HV_LEVEL 0x17, 0" // set HV 0 for H11 | |||
| Send "SCI_HV_LEVEL 0x18, 0" //set HV 0 for H12 | |||
| 2:05 | execute "SCD_CH_MASK_HOT3.STOL" | YS | |
| 2:08 | execute "arc_start.STOL" | YS | |
| 2:10 | execute "daq_start.stol" | YS | |
| 2:12 | GUI display started | JH | |
| 1st data file received | |||
| 2:33 | send command "SCI_SW_CALIB_INTV 300" | YS | |
| to receive spi files every 5 minutes | |||
| send "cal_trig_threshold90.stol" | YS | ||
| to change CAL trigger threshold from 100 to 90 | |||
| 3:03 | SCI_SW_OPMODE 0 | YS | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 3:27 | SCI_SW_OPMODE 0 | YS | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 3:51 | SCI_SW_OPMODE 0 | YS | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 4:20 | data windows was closed | YS | |
| 4:47 | SCI_SW_OPMODE 0 | YS | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 6:06 | SCI_SW_OPMODE 0 | YS | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 6:58 | ESC said they are going to send playback data | YS | |
| 8:58 | connection lost (this happened several times this morning) | ||
| 9:56 | SCI_SW_OPMODE 0 | HP | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 10:25 | SCI_SW_OPMODE 0 | TG | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 10:30 | SCI_SW_OPMODE 0 | HP | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 10:35 | Start ramping hodo MB12 first to 3 kV then 3.8 kV | OG | |
| 10:40 | Reducing trigger threshold from 90 to 77 (through 85 & 80) | OG | |
| 20041220-154807.dat first one with 77 DAC threshold | |||
| 10:57 | Trigger threshold to 75 (cal trig rate ~1 Hz, data rate 25 kbps) | OG | |
| New file starts 20041220-155815.dat with 75 threshold | |||
| 16:42 | SCI_SW_FILE_LIST | ||
| 16:43 | SCI_SW_CUSTOM | ||
| Verified uploaded files were not corrupted. | |||
| 16:49 | SCI_SPARE0 0 | SYZ | |
| Updated flight sw to get cross-calibration events. | |||
| SCI_SW_OPMODE 0 | |||
| Reset DAQ to start the updated flight sw. | |||
| Synchronized CAL hv settings. | |||
| SCI_SPARE1 0 | |||
| which executed SFC scripts for channel masking and setting CAL trigger threshold to 100. | |||
| Executed script scd_ch_mask_hot3.stol. | |||
| Executed script cal_trig_threshold77.stol | |||
| Executed script daq_start.stol | |||
| SCI_SW_HDS_ENABLE_AUSTV 0xFFF, 2, 12 to set delta values | |||
| SCI_SW_HDS_PED_UPDATE_RUN 100 to reset STV with new delta values. | |||
| 16:49 | HV ramp up gradually on Hodo MB12 and MB03 trying for 9 kV | OG | |
| MB11 will be next (forward windows causing delays) | |||
| 16:49 | Cal trigger thresholds to 80 DAC | OG | |
| 19:05 | CAL trigger looks high (10-16Hz). | KC | |
| Buffer (PKT0) goes up to 4% | |||
| Cal trigger thresholds to 85 DAC | KC | ||
| SEND command to make H3 ramping up tp 9kV | YS | ||
| 19:32 | SCI_SW_OPMODE 0 | HP | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 19:33 | SCI_SW_HV_LEVEL 0x0A,3500 | YS | |
| Check the status of HV | |||
| It says H3 area already in 3500 (hkview shows 8.0kV) and H11 are already in 1500 (4.0kV) | |||
| 19:39 | SCI_SW_HV_LEVEL 0x0A,3750 | ||
| SCI_SW_HV_LEVEL 0x12,1750 | |||
| H3 became 8.6kV and H11 became 4.5 KV | |||
| 19:44 | SCI_SW_HV_LEVEL 0x12,2000 (2250) | ||
| 19:57 | Data connection closed because TDRSS is below the horizon | KC | |
| 8:08 | Data resumed | KC | |
| 8:13 | SCI_SW_HV_0x12, 2500 (2750) | KC | |
| Playback started | |||
| 8:32 | SCI_SW_OPMODE 0 | KC | |
| SCI_SNIO_CONTROL 21,0 | |||
| Execute script daq_restart.stol | |||
| 8:33 | SCI_SW_HV_0x12, 3000 | ||
| 8:34 | SCI_SW_HV_0x04, 750 (MB5, Cal ramping) | ||
| 8:36 | MB11 of HDS is now 9.24kv | ||
| 11:41 | Data and command window open | KC | |
| SCI_SW_HV_0x04, 1000 | KC | ||
| 11:50 | Data disconnected | KC | |
| 12:22 | Data connected | KC | |
| 12:32 | SCI_SW_HV 0x04, 1250 | KC | |
| 12:56 | SCI_SW_HV 0x04, 1500 | HP | |
| 1:14 | SCI_SW_HV 0x04, 1750 | KC | |
| 1:22 | SCI_SW_HV 0x04, 2000 | KC | |
| 1:25 | We reached 6.09kv for MB5 of Cal. Successfully and no trip at all so far!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! | KC | |
| 3:15 | Data stopped for 10 minutes due to window time | KC | |
| 4:30 | Data will be stopped for 30 minutes due to window time | KC | |
| 6:40 | Cal trigger thresholds to 82 DAC | KC | |
| 6:46 | Playback started | KC | |
| 7:47 | ESC said will not get data in 30 min | HQ | |
| 8:16 | Start getting data | HQ | |